[figure: 9.7 cm]

LMTO STUDY OF Al_{6}Mn: BAND-STRUCTURE CALCULATIONS AND THE CHARGE DISTRIBUTION

M. Haglauer^{1}, A. Jezierski^{2}, G. Kamieniarz^{1},

1Institute of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań,

Poland

2Institute of Molecular Physics, Polish Academy of Sciences, ul. Smoluchowskiego 17, 60-179 Poznań, Poland

The crystalline Al_{6}Mn has the orthorhombic structure (Cmcm) and includes 28 atoms in a unit cell. The Mn atom is surrounded by ten Al atoms. From the experimental data the electron density distribution (EDD) maps were constructed and the number of valence electrons for each atom was estimated. The charge transfer from Al atoms to an Mn atom on the grounds of both an x-ray diffraction experiment and an analysis by the maximum-entropy method (MEM) was found. Results have shown strong covalent bonds beetween Al and Mn atoms and a strong ionic character of thes bonds [1]. In this comunication, we have studied the electron density distribution of Al_{6}Mn by the ab initio tight binding linearized muffin-tin orbital (TB-LMTO) method. The LMTO calculations have shown that the charge transfer from Al atoms to an Mn atom depends on the Wigner-Saitz radii. This method is used to construct the EDD maps in elementary cell and to calculate the photoemission spectrum.

[1] K. Yamamoto, Y. Matsuo, J. Phys.: Condens. Matter 12, 2359 (2000)

[figure: 13.4 cm]

Subject category :

7. Computations of Electronic Band Structures

Presentation mode : poster

Corresponding author : M. Haglauer

Address for correspondence : Institute of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań, Poland, tel. (+48-61)8273072

Email address : monika@spin.amu.edu.pl